OpenStudy (anonymous):

A little help please?? A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

Should the equation be: \[ |x-0.525|<0.01 \]

OpenStudy (anonymous):

It's writin with the equals sign.

OpenStudy (anonymous):

Then there isn't a minimum or maximum value. There are only two values which it has to be precisely equal to.

OpenStudy (anonymous):

\[ |x-0.525|=0.01 \]Turns into \[ (x-0.525)=0.01\quad\text{and}\quad -(x-0.525)=0.01 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You just have to solve for \(x\) in each equation.

OpenStudy (anonymous):

Thanks I got it now

OpenStudy (anonymous):

Solving for x in each equation above gives you -0.515 and 0.535. Assuming x is the diameter, it can not be negative (obviously). So I'm thinking they need to be within 0.515 and 0.535. Don't forget units if there are any.

OpenStudy (anonymous):

It looks like we are cheating here..this is a quiz question that I'm sure you're going to get caught

OpenStudy (anonymous):

Don't hate